2. If x,y are 2 positive integers and they are derived as follows:
        f(0,y)=y+1.
        f(x,0)=x
        f(x,y)=f(x-1,f(0,y-1))
        f(0,0)=1
          then wats the value of f(3,0),f(1,2) etc nd hw f(1,2) can be derived in terms f(0,0),f(0,2),f(3,0)
              

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

f(0,y) = y+1
f(0,y-1) = y-1+1
f(0,y-1) = y
f(x,0) = x
f(3,0) = 3
f(x,y) = f(x-1,f(0,y-1))
f(x,y) = f(x-1,y)
f(1,0) = f(1-1,0)
f(1,0) = f(0,0) = 1
f(1,2) = f(1-1,2)
f(1,2) = f(0,2)
f(0,2) = 2+1 = 3
f(1,2)=3
f(1,2)=f(0,2)=f(3,0)
f(1,2) = 3f(0,0)